5) Given f(x) = a) (f•g)(2) g(x) = 2x - 4; find the following: c)(g°8X-2)
Answer #1


9 fa) - 1 ____9) = 31- 91) - 27-4 -) (4) Gog) (2) (b) १०१(-२) १०१ = 9(२४-५) = ३(२५-५)-५ - 4K-8-th 909 f09 -f(9(x)) = (२५-47 =